a truck enters a highway at 60mph at the same time a car enters at the same place 6 min later and drives at 69mph , from the time the car entered the highway how long will it take the car to pass the truck?
Answer by Edwin McCravy(20086)


60mph is 1 mile per minute.  So 6 minutes later, when the car enters,
the truck is 6 miles up the road.  The car's approach rate is 9 miles
per hour, found by subtracting their speeds 69-60.  

TIME = DISTANCE/RATE = 6/9 = 2/3 of an hour = 40 minutes.

Edwin